# Unlimited Vector Processing for Wireless Baseband Based on RISC-V Extension

### v1

@@ Abstract
Just as importantly, wireless baseband processing (WBP) demands a more domain-specific architecture [8]. The computation tasks involved in WBP can be broadly categorized into two types. On one hand, tasks such as channel equalization [9] and demodulation [10] are computation-intensive, requiring successive arithmetic operations. On the other hand, tasks like rate-matching [11] involve irregular permutations, for which mainstream architectures offer limited optimization. Moreover, WBP exhibits a high tolerance for quantization error [12], enabling signal processing with fixed-point arithmetic. This allows the large silicon area typically allocated to floating-point units to be repurposed for more domain-specific fixed-point datapaths tailored to WBP, such as complex number computations and saturation logic, which will be discussed in the following sections. The question now is how to design a reasonable instruction set extension for WBP.
@@ II-B Vector Strip-mining
Strip-mining arises as a trade-off among data parallelism, hardware utilization, and software feasibility. However, when applied to long vector processing, it encounters two primary challenges. Firstly, the strip size is constrained by the capacity of vector registers. Although RVV employs power-of-two register group multipliers to mitigate this limitation, the overhead of strip-mining – such as increment/decrement, comparisons, and branching – is not negligible in long vector scenarios. Furthermore, an improper strip size can lead to fragmented memory access patterns. Secondly, vector processing is limited by the number of available vector registers, especially when configuring a bloated LMUL. When multiple operations are bundled within a strip-mining loop, frequent register spilling and filling can occur. These two challenges not only increase memory traffic but also degrade overall throughput, further exacerbating inefficiencies in long vector processing.
@@ III-B1 Vector Register Files
The VRF usage of UVP offers greater flexibility, with RGs dynamically adjusted based on different instructions. RG is determined offline by two factors: the starting register number, , allocated by RA algorithms of the compiler, and the vector data type, defined as , where is the VL specified by the programmer through high-level language. At compile time, RGs are allocated by calculating its range from to . During runtime, the hardware accesses vector registers based on the AVL register specified in the instruction encoding, which allows for cases where the user-defined size exceeds the actual runtime vector size.
@@ III-C1 Matrix Multiplication
As shown in Fig. 3LABEL:sub@fig:matmul, matrix multiplication (matmul) on UVP can be performed in the following steps. ❶ Permutation matrix generation. Two permutation matrices can be generated as follows:
@@ III-C2 Fast Fourier Transform
Fig. 3LABEL:sub@fig:fft illustrates a decimation-in-frequency (DIF) fft kernel with several enhancements. The expanded VRFs allow for the computation of larger-point FFTs. Scatter instructions, combined with permutation vectors and mask flags (marked in red), are employed to re-organize vectors after the butterfly calculation for the next stage of computation. These vectors can either be pre-stored or calculated at runtime, depending on throughput requirements. The instruction count is further optimized by combining the real and imaginary parts of the signal into one single vector. During the calculation, a specialized complex multiplication instruction, uvp_cplxmul, reduces the instruction count from six (four multiplications and two add/subtract instructions) to a single instruction, streamlining the computation and enhancing dataflow efficiency. Moreover, the vshamt register can be pre-configured using the uvp_setcsr instruction to handle data overflow during calculations, which eliminates the need for additional overflow management operations.
@@ IV-C Computation within Lanes
Fix-point division requires careful consideration of two issues: (i) Direct division can lead to underflow, resulting in imprecise number representation and loss of fractional details; and (ii) the quotient may overflow when the dividend is sufficiently small. For a number represented with an -bit integer part and an -bit fractional part, noted as Q, the full-precision calculation of between a divisor of Q and a dividend of Q produces a quotient format of Q. To mitigate underflow, an additional left shift of can be applied to the divisor using vshamt CSR. This shift increases the fractional precision of the quotient at the cost of reducing its integer precision, yielding a quotient format of Q. As illustrated in 4(d), a configurable left shift logic is implemented before the divisor input. To handle overflow, a detection mechanism is implemented using an XOR gate. The gate takes the sign bits of the divisor, dividend, and unsaturated quotient as inputs. Based on the sign of the unsaturated quotient, the output is saturated to its minimum or maximum value accordingly.
@@ V-A Kernel Comparison
We select imatmul and fft from [40] as benchmarks to compare with our design. To align with the execution unit configurations, we modified the element type of the vectors to int16 in each kernel. At the hardware level, the number of lanes in UVP is set to be four times that of Ara, due to the wider SRAMs and the increased DLP of Ara. Clock cycles are measured using Synopsys VCS and QuestaSim, respectively.
Fig. 6 shows the kernel speedup across four lane configurations. matmul (m,n,k) represents an matrix multiplication, and fftN represents a -point FFT. The speedup for matmul ranges from 1.1 to 3.0, with the greatest improvements observed when the matrix dimensions are not powers of two, achieving at least 1.4 speedup. In the proposed matmul kernel example, RGs under different AVL () configurations are more compact compared to the power-of-two RGs. This results in fewer strip-mining loops and switches, as well as register vacancies. For large, power-of-two AVLs, the speedup gain is smaller and primarily results from the reduction in strip-mining loops due to larger strips.

### v2

@@ Abstract
Just as importantly, wireless baseband processing (WBP) demands a more domain-specific architecture [8]. The computation tasks involved in WBP can be broadly categorized into two types. On one hand, tasks such as channel equalization [9] and demodulation [10] are computation-intensive, requiring successive arithmetic operations. On the other hand, tasks like rate-matching [11] involve irregular VLs and permutations, for which mainstream architectures offer limited optimization. Moreover, WBP exhibits a high tolerance for quantization error [12], enabling signal processing with fixed-point arithmetic. This allows the large silicon area typically allocated to floating-point units to be repurposed for more domain-specific fixed-point datapaths tailored to WBP, such as complex number computations and saturation logic, which will be discussed in the following sections. The question now is how to design a reasonable instruction set extension for WBP.
@@ II-A3 Leveraging Scalar Cores for Vector Processing
II-A4 Digital Signal Processors for WBP
Very long instruction word (VLIW) architectures have been widely adopted in digital signal processors (DSP) for their ability to exploit instruction-level parallelism (ILP) through static scheduling. In VLIW processors, multiple operations – including scalar, vector, and even tensor [39] instructions – are bundled and issued simultaneously in a single long instruction word by the compiler [40, 41]. The hardware is greatly simplified compared to superscalar or vector designs, since scheduling, hazard avoidance, and resource allocation are all handled by the compiler. This makes VLIW especially attractive in domain-specific scenarios with predictable workloads and tight power constraints, namely computer vision tasks [42] and WBP [43, 44].
While VLIW DSPs offer high efficiency under regular and static analyzable workloads, they suffer from several limitations. First, the burden of complex scheduling and register allocation falls entirely on the compiler, which demands deep expertise in the target architecture to achieve optimal performance – thereby reducing portability and increasing development effort. Second, static instruction bundling lacks flexibility in handling dynamic workloads or irregular data patterns, often resulting in suboptimal hardware utilization.
@@ II-B Vector Strip-mining
Strip-mining arises as a trade-off among data parallelism, hardware utilization, and software feasibility. However, when applied to long vector processing, it encounters two primary challenges. Firstly, the strip size is constrained by the capacity of vector registers. Although RVV employs power-of-two register group multipliers to mitigate this limitation, the overhead of strip-mining – such as increment/decrement, comparisons, and branching – is not negligible in long vector scenarios. Furthermore, an improper strip size can lead to fragmented memory access patterns. Secondly, vector processing is limited by the number of available vector registers, especially when configuring a bloated LMUL. When multiple operations are bundled within a strip-mining loop, frequent register spilling and filling can occur. These two challenges not only increase memory traffic but also degrade overall throughput, further exacerbating inefficiencies in long vector processing. Compared to conventional vector programming, UVP leverages its support for non-power-of-two RG, effectively alleviating the issues of register under-utilization and excessive strip-mining instructions. Even when the VL is a power of two, UVP still demonstrates advantages due to reduced instruction fetch overhead and fewer pipeline stalls.
@@ III-B1 Vector Register Files
The VRF usage of UVP offers greater flexibility, with RGs dynamically adjusted based on different instructions. RG is determined offline by two factors: the starting register number, , allocated by RA algorithms of the compiler, and the vector data type, defined as , where is the VL specified by the programmer through high-level language. At compile time, RGs are allocated by calculating its range from to . During runtime, the hardware accesses vector registers based on the AVL register specified in the instruction encoding, which allows for cases where the user-defined size exceeds the actual runtime vector size. This design introduces additional complexity in hardware implementation, as UVP requires managing two degrees of freedom ( and ), in contrast to traditional ISAs, which typically rely on only one ().
@@ III-C1 Matrix Multiplication
As shown in Fig. 3, matrix multiplication (matmul) on UVP can be performed in the following steps. ❶ Permutation matrix generation. Two permutation matrices can be generated as follows:
@@ III-C2 Fast Fourier Transform
Fig. 4 illustrates a decimation-in-frequency (DIF) fft kernel with several enhancements. The expanded VRFs allow for the computation of larger-point FFTs. Scatter instructions, combined with permutation vectors and mask flags (marked in red), are employed to re-organize vectors after the butterfly calculation for the next stage of computation. These vectors can either be pre-stored or calculated at runtime, depending on throughput requirements. The instruction count is further optimized by combining the real and imaginary parts of the signal into one single vector. During the calculation, a specialized complex multiplication instruction, uvp_cplxmul, reduces the instruction count from six (four multiplications and two add/subtract instructions) to a single instruction, streamlining the computation and enhancing dataflow efficiency. Moreover, the vshamt register can be pre-configured using the uvp_setcsr instruction to handle data overflow during calculations, which eliminates the need for additional overflow management operations.
@@ IV-C Computation within Lanes
Fix-point division requires careful consideration of two issues: (i) Direct division can lead to underflow, resulting in imprecise number representation and loss of fractional details; and (ii) the quotient may overflow when the dividend is sufficiently small. For a number represented with an -bit integer part and an -bit fractional part, noted as Q, the full-precision calculation of between a divisor of Q and a dividend of Q produces a quotient format of Q. To mitigate underflow, an additional left shift of can be applied to the divisor using vshamt CSR. This shift increases the fractional precision of the quotient at the cost of reducing its integer precision, yielding a quotient format of Q. As illustrated in 5(d), a configurable left shift logic is implemented before the divisor input. To handle overflow, a detection mechanism is implemented using an XOR gate. The gate takes the sign bits of the divisor, dividend, and unsaturated quotient as inputs. Based on the sign of the unsaturated quotient, the output is saturated to its minimum or maximum value accordingly.
@@ V-A Kernel Comparison
We select imatmul, fft, and redsum kernels from [46] as benchmarks to compare with our design. In addition, a minimum mean square error (MMSE) kernel was implemented to compare with [47] in multiple-input, multiple-output (MIMO) systems. To align with the execution unit configurations, we modified the element type of the vectors to int16 in each kernel. At the hardware level, the number of lanes in UVP is set to be four times that of Ara, due to the wider SRAMs and the increased DLP of Ara. Clock cycles are measured using Synopsys VCS and QuestaSim, respectively. Fig. 7 shows the kernel speedup across four lane configurations. matmul (m,n,k) represents an matrix multiplication, fftN represents a -point FFT, and redsumM indicates a reduction summation over elements. For MMSE estimation, we adopt a window size, consistent with [47].
V-A1 Matrix Multiplication
The speedup for matmul ranges from 1.1 to 3.0, with the greatest improvements observed when the matrix dimensions are not powers of two, achieving at least 1.4 speedup. In the proposed matmul kernel example, RGs under different AVL () configurations are more compact compared to the power-of-two RGs. This results in fewer strip-mining loops and switches, as well as register vacancies. An additional source of speedup lies in the loop iteration strategies: Ara’s kernel operates at a complexity of , while the UVP kernel achieves by trading off increased space complexity to store matrix elements. Here, denotes the block size used to configure vsetvli instructions.
V-A2 Fast Fourier Transform
@@
V-A3 Reduction Summation
Reduction operations are widely used in applications such as low-density parity-check (LDPC) decoding and signal correlation, and continue to be the subject of performance optimization efforts [48, 49]. The fourth row of Fig. 7 shows the performance of reduction summation for both power-of-two and non-power-of-two input lengths. Despite having four times the number of lanes compared to Ara – requiring two additional stages of element permutation – UVP consistently achieves at least a 1.2 speedup. This improvement is attributed to less strip-mining and configuration instructions, as well as the elimination of scalar reductions across vector instructions.
V-A4 MMSE Estimation
The complexity of matrix inversion has prompted extensive optimization efforts across platforms [50, 51], with Cholesky decomposition being the most commonly adopted method. The bottom row of Fig. 7 compares the performance of our implementation with that of [47], using a matrix window for a MIMO system. To avoid under-utilization of the VRF, multiple MMSE kernels are executed in parallel under larger lane configurations. UVP still outperforms the Ara implementation, owing to its flexible element reorganization for matrix operations and enhanced support for complex multiplications.
@@ V-C1 Comparison with the State-of-the-Arts
UVP falls short of achieving the maximum frequency due to two primary factors. First, each lane incorporates densely packed arithmetic units which introduce high fan-in logic, often implemented using dense and-or-invert (AOI) cells. This leads to significant local routing congestion. Second, the centralized EXE must maintain datapath connectivity with all surrounding lanes, resulting in placement contention and increased routing complexity.